# Computation method for perfectoid purity and perfectoid BCM-regularity

### v5

@@ Appendix B Test ideal of graded ring
Let $S:=\bigoplus_{m\in\mathbb{Z}_{\geq 0}}S_{m}$ be a Noetherian graded ring such that $S_{0}$ is a field of characteristic $p>0$. We assume $S$ is normal, quasi-Gorenstein and $a(S)<0$ and set $\mathfrak{m}:=S_{>0},d:=\dim{S}$. We further assume $\operatorname{Spec}{S}\backslash\{\mathfrak{m}\}$ is strongly $F$-regular. Then we have
@@
Let $\mathfrak{q}\subseteq S$ be a $\mathfrak{m}$-primary ideal and a rational number $01>t>$ 0. Then there exists a positive integer $e_{0}$, we have $u^{e}(F^{e}_{*}(\mathfrak{q}\cdot\mathfrak{m}^{\lceil t(p^{e}-1)\rceil}))=u^{e}(F^{e}_{*}S)$ for $e\geq e_{0}$.
@@
We may assume $\mathfrak{q}=\mathfrak{m}^{n}$ for some positive integer $n$. We take a rational number $01>t>$ 0. We take $e_{0}$ such that $n+t(p^{e_{0}}-1)\leq p^{e_{0}}+\nu-1$ and $e\geq e_{0}$. Since $\mathfrak{m}^{l}\subseteq S_{\geq l}$ for every positive integer $l$, we have
@@
For every rational number $01>t>$ 0, we have
@@
Finally, we prove the assertion. We take rational numbers $01>t>\varepsilon>$ 0, then we have
@@ Appendix C Proof of Fedder-type criterion
$[\sum_{j=1}^{r}a_{j}f_{j}]\equiv\sum_{j=1}^{r}[a_{j}f_{j}]\ \textrm{mod}\ V(W_{n-1}((f_{1},\ldots,f_{r})))$

### v7

@@ Appendix B Test ideal of graded ring
[…]
@@
Let $\mathfrak{q}\subseteq S$ be a $\mathfrak{m}$-primary ideal and a rational number $1>t>0$. Then there exists a positive integer $e_{0}$, we have $u^{e}(F^{e}_{*}(\mathfrak{q}\cdot\mathfrak{m}^{\lceil t(p^{e}-1)\rceil}))=u^{e}(F^{e}_{*}S)$ for $e\geq e_{0}$.
@@
We may assume $\mathfrak{q}=\mathfrak{m}^{n}$ for some positive integer $n$. We take a rational number $1>t>0$. We take $e_{0}$ such that $n+t(p^{e_{0}}-1)\leq p^{e_{0}}+\nu-1$ and $e\geq e_{0}$. Since $\mathfrak{m}^{l}\subseteq S_{\geq l}$ for every positive integer $l$, we have
@@
For every rational number $1>t>0$, we have
@@
Finally, we prove the assertion. We take rational numbers $1>t>\varepsilon>0$, then we have
@@ Appendix C Proof of Fedder-type criterion
$\sum_{j=1}^{r}a_{j}f_{j}]\equiv\sum_{j=1}^{r}[a_{j}f_{j}]\ \textrm{mod}\ V(W_{n-1}((f_{1},\ldots,f_{r})))$